OpenStudy (anonymous):

Find the given term in each geometric sequence. Hint: Take the 1st number in the pattern x the rate of the pattern to the exponent of (term wanted minus 1). 9th term: 4.5, 9, 18, 36 7th term: 2, 6, 18, 54 6th term: 25, 75, 225, 675 A.) 6075 B.) 1458 C.)1152

jimthompson5910 (jim_thompson5910):

In the sequence 4.5, 9, 18, 36 the common ratio is 2 since 18/9 = 2 or 36/18 = 2

jimthompson5910 (jim_thompson5910):

so r = 2

jimthompson5910 (jim_thompson5910):

the first term is a = 4.5

jimthompson5910 (jim_thompson5910):

the nth term is therefore an = a*r^(n-1) an = 4.5*2^(n-1)

jimthompson5910 (jim_thompson5910):

This formula can now be used to find any term you want In this case, we want the 9th term, so... an = 4.5*2^(n-1) a9 = 4.5*2^(9-1) ... plug in n = 9 a9 = 4.5*2^8 a9 = 4.5*256 a9 = 1152 ------------------------------------------------------- So the ninth term of the sequence 4.5, 9, 18, 36, ... is the number 1152

OpenStudy (anonymous):

an then 7th term: is 1458 an 6th term: is 6075?

jimthompson5910 (jim_thompson5910):

nailed them both

OpenStudy (anonymous):

yay thanks for the help!(:

jimthompson5910 (jim_thompson5910):

you're welcome